OpenStudy (anonymous):

The volume of the cylinder is 1,311.9. Find the height of the cylinder to the nearest tenth of an inch. How do i do this?

OpenStudy (johnweldon1993):

Well the volume of a cylinder is \[\large V = \pi r^2 h\] If we know Volume to be 1311.9 \[\large 1331.9 = \pi r^2h\] Now this as it is is impossible to do since we do not have the radius length... Does the question have that listed?

OpenStudy (anonymous):



OpenStudy (johnweldon1993):

Ah 12... So we have \[\large 1331.9 = 12^2 \pi h\] \[\large 1331.9 = 144 \pi h\] \[\large h = \frac{1331.9}{144 \pi}\]

OpenStudy (anonymous):

so the height would be 9.1in?

OpenStudy (johnweldon1993):

Not quite what I get...

OpenStudy (anonymous):

wait i think i did it wrong that time, is it 2.9?

OpenStudy (johnweldon1993):

^ There we go! :)

OpenStudy (anonymous):

thank you! :)

OpenStudy (johnweldon1993):

Anytime :)